All Systems Operational
Skyline Desktop   Operational
Skyline Mail   Operational
Skyline Voice   Operational
MX Shield Mail Filtering   Operational
SkyRocket Hosting Platform   Operational
DataGuard Online Backup   Operational
Support Centre   Operational
Internet Services Operational
90 days ago
99.94 % uptime
Today
Liverpool PoP (liv-pop-e)   Operational
90 days ago
100.0 % uptime
Today
Liverpool PoP (liv-pop-h)   Operational
90 days ago
99.82 % uptime
Today
Liverpool PoP (liv-pop-h-b)   Operational
90 days ago
99.82 % uptime
Today
Liverpool PoP (liv-pop-a)   Operational
90 days ago
100.0 % uptime
Today
Liverpool PoP (liv-pop-a-m)   Operational
90 days ago
100.0 % uptime
Today
St Helens PoP (sth-pop-a)   Operational
90 days ago
100.0 % uptime
Today
St Helens PoP (sth-pop-b)   Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Scheduled Maintenance
Carrier Notification of Works Dec 7, 00:01-06:01 GMT
There will be a loss of service /redundancy at some point during the timeframe stated above whilst we complete the maintenance work.

This change is to investigate a fix a low light level. All services will be affected for the duration of the investigation
Posted on Nov 26, 16:30 GMT
Past Incidents
Dec 10, 2018

No incidents reported today.

Dec 9, 2018

No incidents reported.

Dec 8, 2018

No incidents reported.

Dec 7, 2018

No incidents reported.

Dec 6, 2018
Resolved - All services are now back to normal, any issues should be reported on the normal support channels.
Dec 6, 11:14 GMT
Monitoring - Servers are now back online, we are now only dealing with specific client access issues. If anything isn't working correctly, please let support know.
Dec 6, 09:55 GMT
Update - Affected services are still coming back online approx albeit quickly now.
Dec 6, 08:27 GMT
Identified - Issue has been identified with storage processor on one of our storage arrays. Service has been moved to the other storage processor while we perform further investigation.
Services which were down are coming back online now on the other storage processor.
Dec 6, 08:04 GMT
Investigating - We are currently investing reports of users not able to access some of our systems.
We will update shortly.
Dec 6, 07:47 GMT
Dec 5, 2018

No incidents reported.

Dec 4, 2018
Completed - All works completed.
Dec 4, 22:42 GMT
Verifying - Verifying works are complete successfully.
Dec 4, 21:45 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Dec 4, 21:01 GMT
Scheduled - We will be performing upgrades to our mail servers between 21:00 and 23:00 this evening during which time reboots maybe required.
Mail access will be intermittent.
Dec 4, 09:44 GMT
Completed - The scheduled maintenance has been completed.
Dec 4, 06:00 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Dec 4, 00:01 GMT
Scheduled - Impact: There will be a loss of service at some point during the timeframe stated above whilst we complete the maintenance work.

The service outage should only last for up to 2 hours; however a 6 hour time frame has been given in the event of any unforeseen circumstances.

We need to let you know about some maintenance work that will be taking place on our network, as detailed below. We will be upgrading the software on our network node at Basildon to fix a known issue and prevent an unplanned outage.
Nov 6, 10:03 GMT
Dec 3, 2018
Resolved - Connection has now been stable for several hours, and no further disruptions are anticipated.
Dec 3, 14:29 GMT
Monitoring - Fix has been applied and we are continuing to monitor the connection over the next few hours to ensure this remains stable.
Dec 3, 10:12 GMT
Identified - Issue has been identified with one of the P2P dishes, a fix is currently being implemented.
Dec 3, 09:15 GMT
Investigating - We are investigating reports of intermittent and slow internet access.
Dec 3, 08:59 GMT
Dec 2, 2018

No incidents reported.

Dec 1, 2018

No incidents reported.

Nov 30, 2018
Resolved - There was a brief interruption to power at the PoP which has now been resolved.
Nov 30, 11:03 GMT
Monitoring - Equipment is now back online, we will advice shortly of any further issues.
Nov 30, 10:37 GMT
Investigating - Investigating loss of service from our PoP on Hope Street, Liverpool. Updates to be provided shortly
Nov 30, 10:26 GMT
Nov 29, 2018

No incidents reported.

Nov 28, 2018
Completed - The scheduled maintenance has been completed.
Nov 28, 18:01 GMT
Update - Scheduled maintenance is still in progress. We will provide updates as necessary.
Nov 28, 12:07 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 28, 06:00 GMT
Scheduled - Our carriers are performing maintenance on the network, information below.

Impact: There will be a loss of service at some point during the timeframe stated above whilst we complete the maintenance work.
The service outage duration is usually between 2-4 hours within the 06:00-18:00 hour working window including 2 hours backout plan activity, however an 12 hours timeframe has been given in the event of any unforeseen circumstances
The planned work is Swap out the Amplifier/s for Liverpool.
Oct 30, 09:46 GMT
Nov 27, 2018
Completed - The scheduled maintenance has been completed.
Nov 27, 06:01 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 27, 00:00 GMT
Scheduled - Impact: There will be a loss of service at some point during the timeframe stated above whilst we complete the maintenance work.

The service outage should only last for up to 2 hours; however a 6 hour time frame has been given in the event of any unforeseen circumstances.

There will be upgrading the software on a network node at Perry Barr to fix a known issue and mitigate an unplanned outage.
Nov 6, 09:57 GMT
Nov 26, 2018
Resolved - This is now rebuild and working correctly.
Nov 26, 16:24 GMT
Update - The search index on one of the affected databases has completed, the other is still rebuilding.
Nov 22, 09:23 GMT
Update - Search is rebuilding, we will advise once this is completed.
Nov 21, 11:01 GMT
Monitoring - Mail databases are now back online, checking on search status and will update again shortly.
Nov 21, 10:10 GMT
Update - We are continuing to work on this issue, we have had to take another database on this server offline for a short time. We will update further however we are working on this with the highest priority.
Nov 21, 09:59 GMT
Identified - Issue with database on this server which we are working on.
Nov 21, 09:07 GMT
Investigating - Investing reports of limited or no access to some mailboxes on mbx-ath01. We will update shortly.
Nov 21, 08:33 GMT